OpenStudy (anonymous):

6. Suppose a parabola has vertex (5, –3) and also passes through the point (6, 1). Write the equation of the parabola in vertex form. (1 point)y = (x – 5)2 – 3 y = 4(x – 5)2 – 3 y = 4(x + 5)2 – 3 y = 4(x – 5)2 + 3

OpenStudy (anonymous):

you know it looks like \[y=a(x-5)^2-3\] but you do not know \(a\) yet to find it, replace \(x\) by \(6\), \(y\) by \(1\) and solve for \(a\) you get \[1=a(6-5)^2-3\] \[1=a-3\] \[a=4\]

OpenStudy (anonymous):

So, what does it mean when a=4?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

OHH i get it thanks :)